I have the circumfernce, & I need 2 find the radius(r)& the diameter(d), What's the formula?
Answer by sdmmadam@yahoo.com(530)

I have the circumfernce, & I need 2 find the radius(r)& the diameter(d), What's the formula?
The formula for the circumference of a circle is
2(Pi)r= 2X(22/7)r where r is the radius of the circle
The value of the circumference is known according to the given data.Let it be V
So we have 2X(22/7)X(r) =V ----(1) where the constant V is known
That is 44r/7 = V
Multiplying by 7
44r = 7V
Dividing by 44
r = (7V/44)
And any diameter of a circle = twice the radius of that circle
Therefore d = 2r = 2X(7V/44) = (7V/22)
Answer: Given the circumference of a circle to be the value V
radius of the circle = (7V/44)
and diameter = (7V/22)